OpenStudy (anonymous):

Story problem! HELP! A parent increases a child's allowance by 15% each year. If the allowance is $3 now, how many years will it take to reach $15? a) Write a equation and find the total amount of years

OpenStudy (anonymous):

starting from 3, next year, it will be A = 3+0.15*3=(1.15)*3 the year afterm A = (1.15)*(1.15)*3 = 3*(1.15)^2 after t years, A = 3*(1.15)^t rearrange for t and solve hint- log

OpenStudy (anonymous):

Wait @electrokid do I solve for t at A = 3*(1.15)^t?

OpenStudy (anonymous):

A= amount t = time thet want an equation for "t" in terms of A

OpenStudy (anonymous):

So I solve for A. I am confused how to start this still... sorry :(

OpenStudy (anonymous):

\[A/3=(1.15)^t\\ \ln({A\over3})=t\ln(1.15)\\ t=\frac{\ln(A)-\ln(3)}{\ln(1.15)} \]

OpenStudy (anonymous):

did you understand how I got A=3*(1.15)^t ?

OpenStudy (anonymous):

now, the question gives you the amount $15. and asks how long will it take for that to happen! so, we solve for "t" in the above equation.

OpenStudy (anonymous):

Yes I understand how you got A=3**1.15)^t... Okay thank you :D